# Faster Diffusion Models via Higher-Order Approximation

### v1

@@ 4.2 Convergence guarantees
[…]
Given that $K$ is a fixed constant and $N\asymp\log T$, the total number of score function evaluations of the proposed algorithm is given by $O(NKT)=O(T\log T)$, which can be as small as $\widetilde{O}\big{(}\frac{d^{1+2/K}}{\varepsilon^{1/K}}\big{)}$ (cf. (18)). In particular, when $K$ grows, one can approximately interpret this number of function evaluations as $\widetilde{O}\big{(}\frac{d^{1+o(1)}}{\varepsilon^{1/K}}\big{)}$, resulting in substantial speed-ups compared to prior theory in Li et al., 2024c ; Li et al., 2024a ; Li and Cai, (2024); Wu et al., (2024). In the noisy case when we only have access to inexact scores, our result guarantees that the TV distance between our output distribution and the target distribution scales proportionally to the two error metrics $\varepsilon_{\mathsf{score}}$ and $\varepsilon_{\mathsf{Jacobi}}.$ It is worth noting that: without assumptions on the smoothness of the score estimates, additional assumptions like the one concerning Jacobian errors cannot be avoided in general, as already illustrated by a counterexample in Li et al., (2023) for the probability flow ODE.
Notably, our approach bears some similarity with other higher-order ODE solvers recently developed in the literature, such as the DPM solver (Lu et al.,, 2022). While remarkable empirical success of such methods have been reported, their theoretical footings remain significantly underdeveloped. Recently, Huang et al., 2024a analyzed the theoretical performance of the $K$-th order Runge-Kutta method for solving the probability flow ODE, for any constant order $K$. While their theory is comparable to ours, it requires higher-order smoothness on the score estimates (more precisely, they required the score estimate’s first $(K+1)$-th derivatives to be bounded). In addition, their iteration complexity also scales linearly in the support size of $p_{\mathsf{data}}$; in contrast, our theory is nearly independent from such a support size as long as it is polynomially bounded. Finally, while we were finalizing our paper, we became aware of a concurrent work Huang et al., (2025), which — under a different set of assumptions — also relaxes the high-order smoothness condition previously imposed in Huang et al., 2024a . Compared with our results, their analysis requires a stronger assumption on the support size, and their total-variation bound depends heavily upon the early-stopping parameter (denoted by $\tau$ therein).
@@ 5.1 Additional notation
Before proceeding to the main proof, we find it helpful to introduce an additional set of notation. Fix $x_{T}\in\mathbb{R}^{d}$, we slightly abuse notation by letting $\{x_{\tau_{t,i}}^{(n)}\}$ denote the deterministic sequence defined by the update rule (14) with $x_{\tau_{T,0}}=x_{T}$. We also define $x_{t}:=x_{\tau_{t},0}$ for all $11\leq t\leq T-$ 1. We let $x_{t}^{\star}$ (resp. $Y_{t}^{\star}$) denote the solution of ODE (9) at $\tau=\tau_{t,0}=1-\overline{\alpha}_{t}$ with the initial condition $x_{\tau_{t+1,0}}^{\star}=x_{t+1}$ (resp. $x_{\tau_{t+1,0}}^{\star}=Y_{t+1}$). For any $\tau\in[\tau_{t,K-1},\tau_{t,0}]=[1-\overline{\alpha}_{t-1},1-\overline{\alpha}_{t}]$, we denote by $x_{\tau}^{\star}$ the solution of ODE (9) at $\tau$ with the initial condition $x_{\tau_{t,0}}^{\star}=x_{t}$. For any $10\leq i\leq K-$ 1, $1\leq t\leq T$, $1\leq n\leq N$, we define
@@ 5.2 Preliminary facts
11 11 $\leq\frac{1-\overline{\alpha}_{t}}{1-\overline{\alpha}_{t-1}}\leq 1+\frac{4c_{1}\log T}{T},\quad~{}2\leq t\leq T$ (23c)
@@
Furthermore, if $-\log p_{\overline{X}_{\tau}}(\lambda x_{1}+(1-\lambda)x_{2})\lesssim d\log T$ for all $\lambda\in[0,1]$, then
@@
The following lemma shows that if $-\log p_{\overline{X}_{\tau}}(x_{\tau}^{\star})$ is not too large, then $-\log p_{\overline{X}_{\tau^{\prime}}}(x_{\tau^{\prime}}^{\star})$ can be well-controlled for all $\tau^{\prime}$ sufficiently close to $\tau$.
@@
Suppose that $-\log p_{\overline{X}_{\widetilde{\tau}}}(x_{\widetilde{\tau}}^{\star})\leq\theta d\log T$ for some $\theta>1$. Then for all $|\tau^{\prime}-\widetilde{\tau}|\leq c_{0}\widetilde{\tau}(1-\widetilde{\tau})$ for some sufficiently small constant $c_{0}>0$, one has
@@ 5.3 Main steps for proving Theorem 1
where $C_{4}>0$ is a sufficiently large constant. Then on the event $\mathcal{E}_{t}$, for any $10\leq i\leq K-$ 1, one has
@@
provided that $T\gtrsim d^{2}\log^{2}T$. Here, $\xi_{t}(\cdot)$ is defined in (20a). Furthermore, the following lemma shows that the log densities $\{-\log p_{X_{\tau_{t,i}}}\big{(}\lambda x_{\tau_{t,i}}^{(n+1)}+(1-\lambda)x_{\tau_{t,i}}^{\star}\big{)}\}_{i,n,\lambda\in[0,1]}$ can be well-controlled as long as $x_{\tau_{t,0}}$ is “typical" and the score estimation errors are not extremely large.
@@
If $C_{10}\frac{\theta_{\tau_{t,0}}(x_{\tau_{t,0}})d\log^{2}T+\sqrt{\theta_{\tau_{t,0}}(x_{\tau_{t,0}})\sum_{i,n}\big{(}\varepsilon_{\mathsf{score},t,i}^{(n)}(x_{\tau_{t,i}}^{(n)})\big{)}^{2}d\log^{3}T}}{T}\leq 1$ for some large enough constant $C_{10}>0$, then for all $10\leq i\leq K-1,0\leq n\leq N-$ 1, one has
@@
=$\quad=\int\left\{\left(\left(1+\sum_{t}\xi_{t}(x_{t})+d^{2}\log^{2}T\bigg{(}\frac{d\log^{2}T}{T}\bigg{)}^{K}\right)q_{T}(x_{T})-p_{T}(x_{T})\right)\operatorname{\mathds{1}}\{x_{1}\in\mathcal{E},x_{T}\in\mathcal{I}_{1}\}\right\}{\rm d}x_{T}$
@@
(48)${{\leq}}4\sum_{t=2}^{T}\mathbb{E}_{Y_{T}\sim p_{T}}\left[\frac{\log T}{T}\left(d\varepsilon_{\mathsf{Jacobi},t}(Y_{t})+\sqrt{d\log T}\varepsilon_{\mathsf{score},t}(Y_{t})\right)\frac{q_{t}(Y_{t})}{p_{t}(Y_{t})}\operatorname{\mathds{1}}\{x_{1}\in\mathcal{E},x_{T}\in\mathcal{I}_{1}\}\right]$
@@
$\mathcal{I}_{2}$ :=$:=\{x_{T}~{}|~{}c_{3}\leq S_{\tau}(x_{T})\leq 2c_{3}\},$ (52)
$\mathcal{I}_{3}$ :=$:=\left\{x_{T}~{}|~{}S_{\tau-1}(x_{T})\leq c_{3},\xi_{\tau}(x_{T})\geq c_{3},\frac{q_{\tau-1}(x_{\tau-1})}{p_{\tau-1}(x_{\tau-1})}\leq 8\frac{q_{\tau}(x_{\tau})}{p_{\tau}(x_{\tau})}\right\},$ (53)
$\mathcal{I}_{4}$ :=$:=\left\{x_{T}~{}|~{}S_{\tau-1}(x_{T})\leq c_{3},\xi_{\tau}(x_{T})\geq c_{3},\frac{q_{\tau-1}(x_{\tau-1})}{p_{\tau-1}(x_{\tau-1})}>8\frac{q_{\tau}(x_{\tau})}{p_{\tau}(x_{\tau})}\right\},$ (54)
@@ 6 Discussion and future directions
In this work, we have proposed an algorithm to speed up data generation in diffusion models. The proposed algorithm exploits $K$-th order ODE approximation combined with a successive refining scheme, which provably yields $\varepsilon$-precision (measured by the TV distance between the output distribution and the target distribution) within $\widetilde{O}\big{(}d^{1+2/K}/\varepsilon^{1/K}\big{)}$ iterations. Our theoretical guarantees hold under fairly mild assumptions on the target distribution, with no restrictions on smoothness and log-concavity, and are capable of accommodating imperfect score estimation. Our findings leave open several interesting questions for future investigation. For instance, while our theory allows the order $K$ to be an arbitrary large constant, it is worth exploring whether $K$ can be allowed to scale with $T$ (e.g., whether $K$ can reach the order of $\log T$), as accomplishing this might potentially lead to further acceleration. Additionally, our method is a deterministic sampler designed to speed up diffusion ODE, and it remains unclear whether a similar iteration complexity can be achieved with an accelerated stochastic sampler (i.e., the SDE counterpart). Furthermore, while this work focuses on the iteration complexity of the sampling phase, how to develop an end-to-end theory that also incorporates the score matching phase forms another direction for future study. Finally, if the target data distribution exhibits intrinsic low dimensionality, it would be interesting to examine whether the sampling process be further expedited.
@@ Appendix A Proof of Lemma 7
We claim that if $-p_{\overline{X}_{\tau}}(x_{\tau}^{\star})\lesssim d\theta\log T$ for $\theta\geq c_{6}$, then for any interger $k\geq 0$, there exists a constant $c_{k}>0$ depending only on $k$ such that
@@
(61)${{\leq}}\frac{1}{3}\max_{0\leq i\leq K}\big{\|}x_{\tau_{t,i}}^{(n)}-x_{\tau_{t,i}}^{\star}\big{\|}_{2}^{2}+C\tau^{\prime 2}\frac{\log^{2}T}{T^{2}}\sum_{0\leq i<K}\big{(}\varepsilon_{\mathsf{score},t,i}^{(n)}(x_{\tau_{t,i}}^{(n)})\big{)}^{2}+C\frac{d\tau^{\prime}\log^{3}T}{T^{2}}\bigg{(}\frac{d\log^{2}T}{T}\bigg{)}^{2K},$
@@
for all $10\leq i\leq k-$ 1, provided that $-\log p_{\tau}(x_{\tau}^{\star})\lesssim d\log T$ for all $\tau_{t,k-1}\leq\tau\leq\tau_{t,0}$. As a consequence, as long as $N\geq C_{0}\log T$ for some large enough constant $C_{0}>0$, one has
@@
$1\lesssim\sqrt{\frac{1-\overline{\alpha}_{t}}{1-\overline{\alpha}_{t+1}}}\leq\left\|\frac{\partial x_{\tau_{t,i}}^{(0)}/\sqrt{1-\tau_{t,i}}}{\partial x_{\tau_{t,0}}/\sqrt{1-\tau_{t,0}}}\right\|=\left\|\frac{\sqrt{1-\tau_{t,0}}}{\sqrt{1-\tau_{t,i}}}I\right\|=\frac{\sqrt{1-\tau_{t,0}}}{\sqrt{1-\tau_{t,i}}}\leq 1$
@@
provided that $-p_{\overline{X}_{\tau_{t,i}}}(x_{\tau_{t,i}}^{(n)})\lesssim d\log T$. Here, the first inequality makes use of Lemma 3 and the second inequality holds due to the induction hypotheses (62a) and (62b) for $n$. Similarly, one has
@@
Then, we prove that (75a) holds for $k=k_{0}+1$. We let $e^{\Delta}=:\sum_{k=0}^{\infty}\frac{\delta^{k}}{k!}w_{k}=\sum_{k=0}^{\infty}\frac{\delta^{k}}{k!}w_{k}(x_{0})$ denote the Taylor expansion of $e^{\Delta}$. Then one can show that $w_{0}=1$ and for all $11\leq k\leq k_{0}+$ 1,
@@
$\mathcal{C}_{0}$ =$=\left\{x_{0}\in\mathbb{R}^{d}:\Big{\|}\frac{x_{\tau}^{\star}}{\sqrt{1-\tau}}-x_{0}\Big{\|}_{2}\leq 5c_{7}\sqrt{\frac{d\theta\tau\log T}{1-\tau}}\right\},$
$\mathcal{C}_{\ell}$ =$=\left\{x_{0}\in\mathbb{R}^{d}:5\cdot 2^{\ell-1}c_{7}\sqrt{\frac{d\tau\log T}{1-\tau}}<\Big{\|}\frac{x_{\tau}^{\star}}{\sqrt{1-\tau}}-x_{0}\Big{\|}_{2}\leq 5\cdot 2^{\ell}c_{7}\sqrt{\frac{d\theta\tau\log T}{1-\tau}}\right\},\quad k\geq 1.$
@@
$\frac{\partial^{k}}{\partial\tau^{k}}\int_{x_{0}}p_{X_{0}\,|\,\overline{X}_{\tau}}(x_{0}\,|\,x_{\tau}^{\star})\exp(\Delta)\big{(}\frac{x_{\tau}^{\star}}{\sqrt{1-\tau}}-x_{0}\big{)}\mathrm{d}x_{0}$ (resp. $\frac{\partial^{k}}{\partial\tau^{k}}\int_{x_{0}}p_{X_{0}\,|\,\overline{X}_{\tau}}(x_{0}\,|\,x_{\tau}^{\star})\exp(\Delta)\mathrm{d}x_{0}$). Then for $10\leq k\leq k_{0}+$ 1,
@@
Here, ()(*)( ∗ ) is valid due to Lemma 2 and (78). Similarly, for all $11\leq k\leq k_{0}+$ 1, one can show that

### v2

@@ 4.1 Assumptions
Remark 2.
By leveraging the techniques in Li and Cai, (2024), one might be able to further relax Assumption 1 to a moment condition of the form
$\mathbb{E}[\|X_{0}\|_{2}^{2}]\leq T^{2c_{R}}$
for an arbitrarily large constant $c_{R}$} without affecting our convergence rates. We do not pursue this refinement in the present work for simplicity.
@@ 4.2 Convergence guarantees
$\text{(iteration complexity)}\qquad\widetilde{O}\bigg{(}\frac{d^{1+2/K}}{\varepsilon^{1/K}}\bigg{)}.$ (18)
Given that $K$ is a fixed constant and $N\asymp\log T$, the total number of score function evaluations of the proposed algorithm is given by $O(NKT)=O(T\log T)$, which can be as small as $\widetilde{O}\big{(}\frac{d^{1+2/K}}{\varepsilon^{1/K}}\big{)}$ (cf. (18)). In particular, when $K$ grows, one can approximately interpret this number of function evaluations as
$\widetilde{O}\bigg{(}\frac{d^{1+o(1)}}{\varepsilon^{1/K}}\bigg{)},$
resulting in substantial speed-ups compared to prior theory in Li et al., 2024c ; Li et al., 2024a ; Li and Cai, (2024); Wu et al., (2024). In particular, our algorithm achieves significant acceleration comparing to the standard discretized probability flow ODE sampler (without the successive-refinement step), which attains $\varepsilon$-accuracy with
$\widetilde{O}(d/\varepsilon)$
score function evaluations (Li et al., 2024c, ). In the noisy case when we only have access to inexact scores, our result guarantees that the TV distance between our output distribution and the target distribution scales proportionally to the two error metrics $\varepsilon_{\mathsf{score}}$ and $\varepsilon_{\mathsf{Jacobi}}.$ It is worth noting that: without assumptions on the smoothness of the score estimates, additional assumptions like the one concerning Jacobian errors cannot be avoided in general, as already illustrated by a counterexample in Li et al., (2023) for the probability flow ODE.
Notably, our approach bears some similarity with other higher-order ODE solvers recently developed in the literature, such as the DPM solver (Lu et al.,, 2022), DPM-Solver++ (Lu et al.,, 2025), and UNiPC (Zhao et al.,, 2023). While remarkable empirical success of such methods have been reported, their theoretical footings remain significantly underdeveloped. Recently, Huang et al., 2024a analyzed the theoretical performance of the $K$-th order Runge-Kutta method for solving the probability flow ODE, for any constant order $K$. While their theory is comparable to ours, it requires higher-order smoothness on the score estimates (more precisely, they required the score estimate’s first $(K+1)$-th derivatives to be bounded). In addition, their iteration complexity also scales linearly in the support size of $p_{\mathsf{data}}$; in contrast, our theory is nearly independent from such a support size as long as it is polynomially bounded. Finally, while we were finalizing our paper, we became aware of a concurrent work Huang et al., (2025), which — under a different set of assumptions — also relaxes the high-order smoothness condition previously imposed in Huang et al., 2024a . Compared with our results, their analysis requires a stronger assumption on the support size, and their total-variation bound depends heavily upon the early-stopping parameter (denoted by $\tau$ therein).
@@ 5.1 Additional notation
Before proceeding to the main proof, we find it helpful to introduce an additional set of notation. Fix $x_{T}\in\mathbb{R}^{d}$, we slightly abuse notation by letting $\{x_{\tau_{t,i}}^{(n)}\}$ denote the deterministic sequence defined by the update rule (14) with $x_{\tau_{T,0}}=x_{T}$. We also define $x_{t}:=x_{\tau_{t},0}$ for all $1\leq t\leq T-1$. We let $x_{t}^{\star}$ (resp. $Y_{t}^{\star}$) denote the solution of ODE (9) at $\tau=\tau_{t,0}=1-\overline{\alpha}_{t}$ with the initial condition $x_{\tau_{t+1,0}}^{\star}=x_{t+1}$ (resp. $x_{\tau_{t+1,0}}^{\star}=Y_{t+1}$). For any $\tau\in[\tau_{t,K-1},\tau_{t,0}]=[1-\overline{\alpha}_{t-1},1-\overline{\alpha}_{t}]$, we denote by $x_{\tau}^{\star}$ the solution of ODE (9) at $\tau$ with the initial condition $x_{\tau_{t,0}}^{\star}=x_{t}$. For any $0\leq i\leq K-1$, $1\leq t\leq T$, $1\leq n\leq N$, we define
@@ 5.2 Preliminary facts
$1$ $\leq\frac{1-\overline{\alpha}_{t}}{1-\overline{\alpha}_{t-1}}\leq 1+\frac{4c_{1}\log T}{T},\quad~2\leq t\leq T$ (23c)
@@
Furthermore, if $\log p_{\overline{X}_{\tau}}(\lambda x_{1}+(1-\lambda)x_{2})\lesssim d\log T$ for all $\lambda\in[0,1]$, then
@@
The following lemma shows that if $\log p_{\overline{X}_{\tau}}(x_{\tau}^{\star})$ is not too large, then $\log p_{\overline{X}_{\tau^{\prime}}}(x_{\tau^{\prime}}^{\star})$ can be well-controlled for all $\tau^{\prime}$ sufficiently close to $\tau$.
@@
[…]
@@ 5.3 Main steps for proving Theorem 1
where $C_{4}>0$ is a sufficiently large constant. Then on the event $\mathcal{E}_{t}$, for any $0\leq i\leq K-1$, one has
@@
provided that $T\gtrsim d^{2}\log^{2}T$. Here, $\xi_{t}(\cdot)$ is defined in (20a).
Remark 4.
It is worth noting that Lemma 7 and (5.3) play a crucial role in deriving the desired convergence rate, and establishing them forms one of the most important components of our proof. To control the approximation errors $\|x_{\tau_{t,K-1}}^{\star}-x_{\tau_{t,K-1}}^{(N)}\|_{2}^{2}$ and $\bigg{\|}\frac{\partial x_{\tau_{t,K-1}}^{\star}/\sqrt{1-\tau_{t,K-1}}}{\partial x_{\tau_{t,0}}^{\star}/\sqrt{1-\tau_{t,0}}}-\frac{\partial x_{\tau_{t,K-1}}^{(N)}/\sqrt{1-\tau_{t,K-1}}}{\partial x_{\tau_{t,0}}/\sqrt{1-\tau_{t,0}}}\bigg{\|}$, we prove sharp bounds on high-order derivatives $\bigg{\|}\frac{\partial^{k}}{\partial\tau^{k}}\frac{s_{\tau}^{\star}(x_{\tau}^{\star})}{(1-\tau)^{3/2}}\bigg{\|}_{2}$ for any fixed $k$, and show that the successive refinements $\{x_{\tau_{t,K-1}}^{(n)}\}_{n\geq 1}$ contract at the desired rates for $n\lesssim\log T$. Detailed steps are deferred to the supplementary material due to space limitations.
Furthermore, the following lemma shows that the log densities $\{-\log p_{X_{\tau_{t,i}}}\big{(}\lambda x_{\tau_{t,i}}^{(n+1)}+(1-\lambda)x_{\tau_{t,i}}^{\star}\big{)}\}_{i,n,\lambda\in[0,1]}$ can be well-controlled as long as $x_{\tau_{t,0}}$ is “typical" and the score estimation errors are not extremely large.
@@
If $C_{10}\frac{\theta_{\tau_{t,0}}(x_{\tau_{t,0}})d\log^{2}T+\sqrt{\theta_{\tau_{t,0}}(x_{\tau_{t,0}})\sum_{i,n}\big{(}\varepsilon_{\mathsf{score},t,i}^{(n)}(x_{\tau_{t,i}}^{(n)})\big{)}^{2}d\log^{3}T}}{T}\leq 1$ for some large enough constant $C_{10}>0$, then for all $0\leq i\leq K-1,0\leq n\leq N-1$, one has
@@
$\quad=\int\left\{\left(\left(1+\sum_{t}\xi_{t}(x_{t})+d^{2}\log^{2}T\bigg{(}\frac{d\log^{2}T}{T}\bigg{)}^{K}\right)q_{T}(x_{T})-p_{T}(x_{T})\right)\operatorname{\mathds{1}}\{x_{1}\in\mathcal{E},x_{T}\in\mathcal{I}_{1}\}\right\}{\rm d}x_{T}$
@@
[…]
@@
$\mathcal{I}_{2}$ $:=\{x_{T}~|~c_{3}\leq S_{\tau}(x_{T})\leq 2c_{3}\},$ (52)
$\mathcal{I}_{3}$ $:=\left\{x_{T}~|~S_{\tau-1}(x_{T})\leq c_{3},\xi_{\tau}(x_{T})\geq c_{3},\frac{q_{\tau-1}(x_{\tau-1})}{p_{\tau-1}(x_{\tau-1})}\leq 8\frac{q_{\tau}(x_{\tau})}{p_{\tau}(x_{\tau})}\right\},$ (53)
$\mathcal{I}_{4}$ $:=\left\{x_{T}~|~S_{\tau-1}(x_{T})\leq c_{3},\xi_{\tau}(x_{T})\geq c_{3},\frac{q_{\tau-1}(x_{\tau-1})}{p_{\tau-1}(x_{\tau-1})}>8\frac{q_{\tau}(x_{\tau})}{p_{\tau}(x_{\tau})}\right\},$ (54)
@@ 6 Discussion and future directions
In this work, we have proposed an algorithm to speed up data generation in diffusion models, and have contributed towards a new theoretical framework for understanding the efficacy of higher-order ODE-based sampling methods. The proposed algorithm exploits $K$-th order ODE approximation combined with a successive refining scheme, which provably yields $\varepsilon$-precision (measured by the TV distance between the output distribution and the target distribution) within $\widetilde{O}\big{(}d^{1+2/K}/\varepsilon^{1/K}\big{)}$ iterations. Our theoretical guarantees hold under fairly mild assumptions on the target distribution, with no restrictions on smoothness and log-concavity, and are capable of accommodating imperfect score estimation.
Our findings leave open several interesting questions for future investigation. For instance, while our theory allows the order $K$ to be an arbitrary large constant, it is worth exploring whether $K$ can be allowed to scale with $T$ (e.g., whether $K$ can reach the order of $\log T$), as accomplishing this might potentially lead to further acceleration. Additionally, our method is a deterministic sampler designed to speed up diffusion ODE, and it remains unclear whether a similar iteration complexity can be achieved with an accelerated stochastic sampler (i.e., the SDE counterpart). Furthermore, while this work has focused on the iteration complexity of the sampling phase, how to develop an end-to-end theory that also incorporates the score matching phase forms another direction for future study. Finally, when the target data distribution exhibits intrinsic low dimensionality, it would be interesting to examine whether the sampling process is further expedited.
Acknowledgments
G. Li is supported in part by the Chinese University of Hong Kong Direct Grant for Research and the Hong Kong Research Grants Council ECS 2191363. Y. Wei is supported in part by the NSF grants CCF-2106778, CCF-2418156 and CAREER award DMS-2143215. Y. Chen is supported in part by the Alfred P. Sloan Research Fellowship, the ONR grants N00014-22-1-2354 and N00014-25-1-2344, the NSF grants 2221009 and 2218773, the Wharton AI & Analytics Initiative’s AI Research Fund, and the Amazon Research Award.
@@ Appendix A Proof of Lemma 7
We claim that if $p_{\overline{X}_{\tau}}(x_{\tau}^{\star})\lesssim d\theta\log T$ for $\theta\geq c_{6}$, then for any interger $k\geq 0$, there exists a constant $c_{k}>0$ depending only on $k$ such that
@@
[…]
@@
for all $0\leq i\leq k-1$, provided that $\log p_{\tau}(x_{\tau}^{\star})\lesssim d\log T$ for all $\tau_{t,k-1}\leq\tau\leq\tau_{t,0}$. As a consequence, as long as $N\geq C_{0}\log T$ for some large enough constant $C_{0}>0$, one has
@@
$1\lesssim\sqrt{\frac{1-\overline{\alpha}_{t}}{1-\overline{\alpha}_{t+1}}}\leq\left\|\frac{\partial x_{\tau_{t,i}}^{(0)}/\sqrt{1-\tau_{t,i}}}{\partial x_{\tau_{t,0}}/\sqrt{1-\tau_{t,0}}}\right\|=\left\|\frac{\sqrt{1-\tau_{t,0}}}{\sqrt{1-\tau_{t,i}}}I\right\|=\frac{\sqrt{1-\tau_{t,0}}}{\sqrt{1-\tau_{t,i}}}\leq 1.$
@@
provided that $p_{\overline{X}_{\tau_{t,i}}}(x_{\tau_{t,i}}^{(n)})\lesssim d\log T$. Here, the first inequality makes use of Lemma 3 and the second inequality holds due to the induction hypotheses (62a) and (62b) for $n$. Similarly, one has
@@
Then, we prove that (75a) holds for $k=k_{0}+1$. We let $e^{\Delta}=:\sum_{k=0}^{\infty}\frac{\delta^{k}}{k!}w_{k}=\sum_{k=0}^{\infty}\frac{\delta^{k}}{k!}w_{k}(x_{0})$ denote the Taylor expansion of $e^{\Delta}$. Then one can show that $w_{0}=1$ and for all $1\leq k\leq k_{0}+1$,
@@
$\mathcal{C}_{0}$ $=\left\{x_{0}\in\mathbb{R}^{d}:\Big{\|}\frac{x_{\tau}^{\star}}{\sqrt{1-\tau}}-x_{0}\Big{\|}_{2}\leq 5c_{7}\sqrt{\frac{d\theta\tau\log T}{1-\tau}}\right\},$
$\mathcal{C}_{\ell}$ $=\left\{x_{0}\in\mathbb{R}^{d}:5\cdot 2^{\ell-1}c_{7}\sqrt{\frac{d\tau\log T}{1-\tau}}<\Big{\|}\frac{x_{\tau}^{\star}}{\sqrt{1-\tau}}-x_{0}\Big{\|}_{2}\leq 5\cdot 2^{\ell}c_{7}\sqrt{\frac{d\theta\tau\log T}{1-\tau}}\right\},\quad k\geq 1.$
@@
$\frac{\partial^{k}}{\partial\tau^{k}}\int_{x_{0}}p_{X_{0}\,|\,\overline{X}_{\tau}}(x_{0}\,|\,x_{\tau}^{\star})\exp(\Delta)\big{(}\frac{x_{\tau}^{\star}}{\sqrt{1-\tau}}-x_{0}\big{)}\mathrm{d}x_{0}$ (resp. $\frac{\partial^{k}}{\partial\tau^{k}}\int_{x_{0}}p_{X_{0}\,|\,\overline{X}_{\tau}}(x_{0}\,|\,x_{\tau}^{\star})\exp(\Delta)\mathrm{d}x_{0}$). Then for $0\leq k\leq k_{0}+1$,
@@
Here, ()(*) is valid due to Lemma 2 and (78). Similarly, for all $1\leq k\leq k_{0}+1$, one can show that